OpenStudy (anonymous):

The distance from carrie's house to the park was 3/4 of a mile. Carrie walked 2/3 of the way before she stopped for a drink. How far had Carrie walked?

OpenStudy (anonymous):

all you have to find is what 2/3 of 3/4 is.

OpenStudy (anonymous):

I know.

OpenStudy (anonymous):

How can I?

OpenStudy (anonymous):

i think you divide im pretty sure but i may be wrong

OpenStudy (anonymous):

can someone please reply to mine :(

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so its \[\frac{ 18 }{ 4 }\]

OpenStudy (anonymous):

which is 4 1/2 which doesnt sound right hold on let me think

OpenStudy (anonymous):

yeah ill post the 4 choices

OpenStudy (texaschic101):

2/3 of 3/4 = 2/3 x 3/4 --- " of " means multiply

OpenStudy (anonymous):

i think the answer is 1/2 because of means multiply and 2/3 x 3/4 + 1/2

OpenStudy (anonymous):

i did it the long way in my head though

OpenStudy (anonymous):

here are the choices A. 1/2 mile B. 5/7 mile C. 1/3 mile D. 1/4 mile

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{4}\cdot \cfrac{2}{3}\)

OpenStudy (anonymous):

@jdoe0001 can u help me plz

OpenStudy (anonymous):

its choice a

OpenStudy (jdoe0001):